OpenStudy (anonymous):

Determine the solution of the system. Y = 5x - 9 Y = 2x + 6

OpenStudy (texaschic101):

5x - 9 = 2x + 6 can you solve for x now ??

syed98 (syedmohammed98):

Well u can use substitution or elimination as well :)

OpenStudy (anonymous):

I still need help figure that out i suck at this

OpenStudy (texaschic101):

5x - 9 = 2x + 6 --- subtract 2x from both sides 5x - 2x - 9 = 2x - 2x + 6 -- combine like terms 3x - 9 = 6 --- add 9 to both sides 3x - 9 + 9 = 6 + 9 -- combine like terms 3x = 15 -- divide both sides by 3 x = 5 now sub 5 in for x into either of the equations to find y y = 5x - 9 y = 5(5) - 9 y = 25 - 9 y = 16 check.. y = 2x + 6 16 = 2(5) + 6 16 = 10 + 6 16 = 16 (correct) any questions ? I tried to leave in all the steps so you could understand.

OpenStudy (anonymous):

Thank you that helps alot!

OpenStudy (texaschic101):

glad to help :)